How many liters of a 20% alcohol solution must be added to 80 liters of a 50% alcohol solution to form a 45% solution?
Also,
How many pounds of coffee A at $3/pound should be mixed with 2.5 pounds of coffee B at $4.20/pound to form a mixture selling for $3.75/pound?

Answer by stanbon(75887)

How many liters of a 20% alcohol solution must be added to 80 liters of a 50% alcohol solution to form a 45% solution?
------------
Let amount of 20% solution be "x":
EQUATION:
alcohol: 0.20x + 0.50*80 = 0.45(x+80)
20x + 4000 = 45x + 3600
25x = 400
x = 16 liters (amount of 20% solution that must be added)
---------------------------------

Also,
How many pounds of coffee A at $3/pound should be mixed with 2.5 pounds of coffee B at $4.20/pound to form a mixture selling for $3.75/pound?
EQUATION:
value: 3A + 4.2*2.5 = 3.75*(A + 2.5)
3A + 10.5 = 3.75A + 9.375
0.75A = 1.125
A = 1.5 lbs. (Amount of coffee that must be added)
================
Cheers,
Stan H.